You are monitoring length of corn that your sister is selling at the farmer’s market. You...

  1. You are monitoring length of corn that your sister is selling at the farmer’s market. You have following data over 4 days. Draw 3 sigma level control charts (X-bar and R-chart). Calculate and show values for UCL, CL, and LCL using the table given below. Discuss your graphs. Is everything okay?

day 1

day 2

day 3

day 4

15

13

10

8

13

12

9

7

15

11

11

9

           

Factors for 3 Sigma level

Sample Control Limit Factor          Control Limit Factor

Size          for Sample Mean               for Sample Range

n            A2                       D3                        D4

2            1.880                   0                          3.267

3            1.023                   0                          2.575

4            0.729                   0                          2.282

5            0.577                   0                          2.116

10          0.308                   0.223                   1.777

15          0.223                   0.348                   1.652

20          0.180                   0.414                   1.586

25          0.153                   0.459                   1.541

I need step by step on how to create the graphs. Thank you
Answer #1

day 1

day 2

day 3

day 4

Range

15

13

10

8

15-8=7

13

12

9

7

13-7=6

15

11

11

9

15-9=6

X double bar=133/12=11.08

R bar= (7+6+6)/3=6.33

For n=4, A2= 0.729  , D3=0, D4= 2.282

For X bar chart

CL= 133/12=11.08

UCL= X double bar+A2*R bar =11.08+0.729*6.33=15.69

LCL= X double bar-A2*R bar =11.08-0.729*6.33=6.47

Chart is as below=

For R Chart

CL=19/3=6.33

UCL= D4*R bar=2.282*6.33=14.45

LCL= D3*R bar=0*6.33=0

Chart is given below

As all the points of X bar and R charts are within the UCL and LCL values, thus the process is in control